### v1

# Theoretical Foundations of the Deep Copula Classifier: A Generative Approach to Modeling Dependent Features

Abstract
Traditional classifiers often assume feature independence or rely on overly simplistic relationships, leading to poor performance in settings where real-world dependencies matter. We introduce the Deep Copula Classifier (DCC), a generative model that separates the learning of each feature’s marginal distribution from the modeling of their joint dependence structure via neural network-parameterized copulas. For each class, lightweight neural networks are used to flexibly and adaptively capture feature interactions, making DCC particularly effective when classification is driven by complex dependencies. We establish that DCC converges to the Bayes-optimal classifier under standard conditions and provide explicit convergence rates of $O(n^{-r/(2r+d)})$ for $r$-smooth copula densities. Beyond theoretical guarantees, we outline several practical extensions, including high-dimensional scalability through vine and factor copula architectures, semi-supervised learning via entropy regularization, and online adaptation using streaming gradient methods. By unifying statistical rigor with the representational power of neural networks, DCC offers a mathematically grounded and interpretable framework for dependency-aware classification.
Keywords:Classification, Copula Theory, Deep Generative Models, Feature Dependencies, Bayes Consistency, Neural Copulas
MSC 2020 Classification: 62H30, 68T07, 62C12, 62G05
1 Introduction
Classification, that is the task of assigning labels based on features, is a cornerstone of machine learning. Classical methods like logistic regression [8], support vector machines [5], and naive Bayes [19] succeed in many scenarios but rely on restrictive assumptions: naive Bayes presumes feature independence, linear discriminant analysis requires Gaussian features with shared covariance, and even neural networks encode feature interactions opaquely through weight matrices rather than explicit probabilistic modeling.
Real-world features, however, interact through complex nonlinear dependencies. Financial indicators exhibit extreme co-movements invisible to correlation [7], while medical biomarkers gain diagnostic power only in specific combinations. Ignoring these interdependencies degrades classifier performance. Copulas elegantly solve this by decoupling marginal distributions from dependence structures [27]. Traditional parametric copulas (Gaussian, Clayton, Gumbel) and newer Archimedean variants [1] capture basic patterns but falter in high dimensions or with intricate dependencies. Neural copulas [30, 32, 11] overcome these limitations through adaptive deep learning.
We bridge this gap with the Deep Copula Classifier (DCC), which for each class models individual features via their marginal distributions and learns feature dependencies through neural network-parameterized copulasThis generative approach explicitly quantifies how features co-vary, enabling classifiers that harness dependency patterns. Our theoretical contributions are as follows:
  1. 1.
    Novel Framework: Integration of deep copulas into a class-conditional generative classifier
  2. 2.
    Consistency Proof: DCC converges to Bayes-optimality under standard conditions
  3. 3.
    Rate Quantification: $O(n^{-r/(2r+d)})$ excess risk for $r$-smooth copulas
  4. 4.
    Interpretability: Direct inspection of learned copula dependencies
  5. 5.
    Extensions Roadmap: High-dimensional, semi-supervised, and streaming adaptations
DCC combines copula theory’s rigor with neural networks’ flexibility, offering a principled path for dependency-aware classification. This foundation supports both theoretical analysis and practical implementation.
2 Related Work
2.1 Copula Theory and Foundations
The copula framework, formalized by Sklar’s theorem [27], enables the separation of marginal distributions from joint dependence structures. Modern treatments, such as [22] establish copulas as foundational tools in multivariate analysis. Significant progress has been made in several directions. Nonparametric estimation techniques, including kernel-based copula density estimators [23] and empirical bayes method [18], have enabled flexible inference without strong distributional assumptions. In high dimensions, vine copulas [21] offer a scalable strategy by factorizing complex dependencies into cascades of bivariate copulas. Additionally, the integration of copulas into generative modeling frameworks, such as copula-based variational inference [28] and optimal transport-based constructions [4], has expanded their utility in probabilistic learning.
2.2 Generative Classification Paradigms
Traditional generative classifiers, such as naive Bayes [19], prioritize tractability over accurate modeling of feature dependence, often assuming conditional independence. In contrast, modern deep generative approaches address this limitation through implicit and explicit mechanisms. Models like variational autoencoders (VAEs) [14] and generative adversarial networks (GANs) [10] capture feature interactions within nonlinear latent spaces, providing powerful implicit dependence structures. Others, such as neural graphical model hybrids [26], incorporate explicit conditional independence assumptions to combine interpretability with deep learning. The Deep Copula Classifier (DCC) diverges from both approaches by directly parameterizing feature dependencies through copulas while preserving full generative semantics, a principled middle ground between rigid parametric assumptions and black-box neural inference.
2.3 Neural Copula Architectures
Recent advances in neural dependence modeling provide key building blocks for flexible and expressive copula-based models. Universal parameterization has been achieved using deep Archimedean copulas [16] and transformer-based copula processes [32], allowing for rich representations of complex multivariate dependencies. For adaptive estimation, nonparametric vine copulas augmented with neural pair-copula blocks have shown promise in capturing localized interactions in high dimensions [21]. Additionally, hybrid learning paradigms such as neural copula processes [11] effectively blend generative and discriminative objectives, offering a unified framework for robust prediction and inference.
Our work extends these foundations by developing:
  1. 1.
    Class-specific copula networks with theoretical consistency guarantees
  2. 2.
    Explicit finite-sample convergence rates under smoothness conditions
  3. 3.
    Architectural constraints ensuring valid density outputs
    This bridges the gap between empirical neural copula methods and rigorous statistical learning theory.
3 Theoretical Framework
3.1 Copula Basics
We begin by recalling how copulas enable separation between marginal distributions and joint dependence patterns.
Definition 1 (Copula).
A $d$-dimensional copula is a function $C:[0,1]^{d}\to[0,1]$ satisfying:
  1. 1.
    Groundedness: If any $u_{i}=0$, then $C(u_{1},\dots,u_{d})=0$.
  2. 2.
    Uniform margins: For each $i$,
    $C(1,\dots,1,u_{i},1,\dots,1)=u_{i}.$
  3. 3.
    $d$-increasing: $C$ assigns non-negative volume to every axis-aligned box in $[0,1]^{d}$.
Theorem 1 (Sklar’s Theorem [27]).
Let $F$ be any joint distribution on $(X_{1},\dots,X_{d})$ with continuous marginals $F_{1},\dots,F_{d}$. Then there exists a unique copula $C$ such that
$F(x_{1},\dots,x_{d})=C(F_{1}(x_{1}),\dots,F_{d}(x_{d})).$
Conversely, any copula $C$ combined with valid marginals $F_{i}$ produces a joint distribution through this relationship.
When $C$ is differentiable, we obtain its density via:
$c(u_{1},\dots,u_{d})=\frac{\partial^{d}}{\partial u_{1}\cdots\partial u_{d}}C(u_{1},\dots,u_{d}).$
3.2 Deep Copula Classifier Definition
Our classifier models each class’s joint distribution through neural copula parameterization.
Definition 2 (Deep Copula Classifier).
Let $X=(X_{1},\dots,X_{d})$ be a feature vector and $Y\in\{1,\dots,K\}$ a discrete label. For each class $y$, we decompose the conditional density as:
$P(X\mid Y=y)=c_{y}(F_{1|y}(X_{1}),\dots,F_{d|y}(X_{d}))\cdot\prod_{i=1}^{d}f_{i|y}(X_{i}),$
where:
  • $F_{i|y}$ and $f_{i|y}$ are the class-conditional marginal CDF and PDF;
  • $c_{y}$ is a copula density implemented through a neural network $NN_{y}(\cdot;\theta_{y})$ with softplus output activation and implicit normalization.
The transformation $u_{i}=F_{i|y}(X_{i})$ maps features to the unit hypercube, where $NN_{y}$ learns dependence patterns. While practical implementations require explicit density normalization, our theoretical analysis assumes the network can approximate any smooth copula density through Hornik’s universal approximation theorem [12].
3.3 Consistency and Convergence
We establish fundamental guarantees under the following conditions:
Assumption 1 (Regularity Conditions).
Assume:
  1. 1.
    True class-conditional distributions permit copula factorizations.
  2. 2.
    Marginals $F_{i|y}$ are continuous and strictly increasing.
  3. 3.
    Neural networks have width/depth growing polynomially with $n$ to satisfy Hornik’s approximation requirements.
  4. 4.
    Training samples $(X^{(i)},Y^{(i)})$ are i.i.d.
  5. 5.
    Marginal estimators $\hat{F}_{i|y}$ converge at rate $O(n^{-\alpha})$ with $\alpha>\frac{r}{2r+d}$.
Theorem 2 (Consistency).
Under Assumption 1, as $n\to\infty$ with network capacity growing appropriately, the DCC achieves Bayes-optimal classification:
$\lim_{n\to\infty}P(\hat{Y}\neq Y\mid X)=P^{*}(Y\neq Y^{*}\mid X).$
Theorem 3 (Convergence Rate).
If true copula densities belong to a Sobolev space $W^{r}_{2}([0,1]^{d})$ of order $r$, the DCC’s excess risk decays as:
$E(n)=O\left(n^{-\frac{r}{2r+d}}+n^{-\alpha}\right)=O\left(n^{-\frac{r}{2r+d}}\right)$
neglecting logarithmic factors.
Detailed proofs appear in Appendix ( Appendix A: Detailed Proofs).
4 Algorithm and Implementation
4.1 Training Procedure
The DCC training process decomposes into marginal estimation and dependence learning, as formalized in Algorithm 1.
Algorithm 1 Theoretical Training Protocol for DCC
0:  Labeled sample $\{(x^{(i)},y^{(i)})\}_{i=1}^{n}\stackrel{{\scriptstyle iid}}{{\sim}}P_{X,Y}$
0:  Estimated marginals $\{\hat{F}_{j|y},\hat{f}_{j|y}\}$ and parameters $\{\hat{\theta}_{y}\}_{y=1}^{K}$
1:  Phase 1: Consistent Marginal Estimation
2:  for each class $y\in\{1,\dots,K\}$ do
3:     for each feature $j\in\{1,\dots,d\}$ do
4:        Estimate $\hat{F}_{j|y}$ via smoothed empirical CDF with bandwidth $h=o(n^{-1/2})$
5:        Derive $\hat{f}_{j|y}=\frac{d}{dx}\hat{F}_{j|y}$ using kernel density estimation
6:     end for
7:  end for
8:  Phase 2: Neural Copula Learning
9:  for each class $y\in\{1,\dots,K\}$ do
10:     Transform data: $u_{j}^{(i)}=\hat{F}_{j|y}(x_{j}^{(i)})$ for all $i$ such that $y^{(i)}=y$
11:     Define $NN_{y}$ with architecture from Section 4.3
12:     Solve $\hat{\theta}_{y}=\arg\max_{\theta}\frac{1}{n_{y}}\sum_{i:y^{(i)}=y}\log NN_{y}(u^{(i)};\theta)$
13:     where $n_{y}=|\{i:y^{(i)}=y\}|$
14:  end for
4.2 Prediction Procedure
Classification follows the Bayes decision rule under estimated distributions (Algorithm 2).
Algorithm 2 Theoretical Prediction Mechanism
0:  Test point $x^{*}$, trained DCC model
0:  Predicted class $\hat{y}$
1:  for each class $y\in\{1,\dots,K\}$ do
2:     Transform: $u_{j}^{*}=\hat{F}_{j|y}(x_{j}^{*})$ for $j=1,\dots,d$
3:     Evaluate: $\hat{c}_{y}(u^{*})=NN_{y}(u^{*};\hat{\theta}_{y})$
4:     Compute: $L_{y}=\hat{c}_{y}(u^{*})\cdot\prod_{j=1}^{d}\hat{f}_{j|y}(x_{j}^{*})$
5:     Incorporate prior: $\tilde{L}_{y}=L_{y}\cdot\hat{\pi}_{y}$, where $\hat{\pi}_{y}=\frac{n_{y}}{n}$
6:  end for
7:  Return $\hat{y}=\arg\max_{y\in\{1,\dots,K\}}\tilde{L}_{y}$
4.3 Neural Network Architecture
The copula networks $NN_{y}$ require architectural constraints to ensure theoretical validity:
[…]
  • Hidden Layers: Depth $L=\lceil\log_{2}n\rceil$ with width $W=O(n^{d/(2r+d)})$ for $r$-smooth copulas
  • Activation: ReLU for hidden layers, Softplus output activation for positivity
  • Regularization: Implicit via spectral normalization [20] to control Rademacher complexity
Note: The integral constraint is theoretically achievable through Barron space approximation [2] given sufficient network capacity. Practical implementations would require normalization techniques, but our theoretical analysis assumes exact integration.
5 Conclusion and Future Work
Our theoretical analysis establishes the Deep Copula Classifier (DCC) as a Bayes-consistent method that explicitly models feature dependencies through neural copula parameterization. By separating marginal learning from dependence estimation, DCC provides an interpretable framework where practitioners can inspect both individual feature behaviors (through $\hat{F}_{j|y}$) and interaction patterns (via $NN_{y}$). The $O(n^{-r/(2r+d)})$ convergence rate for Sobolev-class copulas demonstrates adaptivity to intrinsic data complexity.
The following avenues for theoretical extension emerge:
High-Dimensional Scaling The curse of dimensionality in $[0,1]^{d}$ copula estimation poses significant challenges. To address this, one promising direction involves vine copula factorizations [21], which break down high-dimensional dependencies into tree-structured graphs that are easier to estimate. Additionally, low-rank tensor decompositions of the weight matrices within $NN_{y}$ may reduce parameter complexity while preserving expressive power. Another line of work would analyze the tradeoff between sparsity and sample size in regimes where $d\gg n$, providing theoretical insights into when DCC remains statistically reliable.
Semi-Supervised Learning Given its generative nature, DCC naturally lends itself to semi-supervised extensions under the cluster assumption [3]. This opens the door to convergence analyses for entropy-regularized objectives that leverage both labeled and unlabeled data. In particular, DCC may benefit when marginal distributions $P(X)$ carry information about class boundaries, leading to provable gains in label efficiency. Furthermore, a rigorous comparison between DCC and discriminative semi-supervised methods could clarify the advantages of modeling dependencies explicitly.
Online Adaptation In streaming data settings, DCC can be extended to incorporate online learning mechanisms. One direction is to establish regret bounds for sequential updates to the copula density via mirror descent [13]. Beyond regret analysis, theoretical guarantees will be needed to understand how DCC behaves under class-incremental scenarios, especially regarding catastrophic forgetting. Additionally, DCC provides a unique opportunity to model and detect concept drift through shifts in learned dependence structures over time.
Empirical-Theoretical Synergy Although our focus has been theoretical, empirical observations raise questions that demand mathematical exploration. For example, DCC may exhibit phase transitions in classification performance as the strength of feature dependencies varies, phenomena that could be characterized through theoretical thresholds. Another promising direction involves deriving information-theoretic lower bounds for copula-based classification, which would help identify when dependency modeling provides fundamental advantages over independence-based models.
The DCC framework opens new questions about balancing parametric and nonparametric components in hybrid models. By maintaining explicit probability models rather than black-box representations, it enables precise characterization of when and how feature dependencies impact classification, a crucial step toward explainable AI for high-stakes decisions.
Statements and Declarations
Funding
The authors did not receive any funding for this research.
Competing Interests
The authors have no relevant financial or non-financial interests to disclose.
Appendix A: Detailed Proofs
A.1 Proof of Theorem 2 (Consistency)
We demonstrate convergence to Bayes optimality through three interlocking arguments:
Step 1: Marginal Convergence
Under Assumption 1(point 5), the marginal estimators satisfy:
$\sup_{x}\left|\hat{F}_{i|y}(x)-F_{i|y}(x)\right|=O_{p}(n^{-\alpha})$
This follows from standard nonparametric estimation theory [31, Theorem 2.1] for smoothed CDF estimators (e.g., boundary-corrected kernel methods). The pseudo-observations $\hat{u}_{i}=\hat{F}_{i|y}(X_{i})$ thus satisfy:
$\|\hat{u}_{i}-u_{i}\|_{\infty}=O_{p}(n^{-\alpha})$
Step 2: Neural Copula Approximation
By Assumption 1(point 3) and Hornik’s theorem [12], networks with width/depth polynomial in $n$ can approximate Sobolev-class copula densities. For each class $y$, there exist parameters $\theta_{y}^{*}$ such that:
$\sup_{u\in[0,1]^{d}}\left|NN_{y}(u;\theta_{y}^{*})-c_{y}(u)\right|=O\left(n^{-\frac{r}{2r+d}}\right)$
Maximum likelihood estimation over this class yields $\hat{\theta}_{y}$ satisfying:
$\mathbb{E}_{u}\left[\left|NN_{y}(u;\hat{\theta}_{y})-c_{y}(u)\right|^{2}\right]=O_{p}\left(n^{-\frac{2r}{2r+d}}+n^{-1}\right)$
via Rademacher complexity bounds [15].
Step 3: Decision Rule Stability
The composite density estimate satisfies:
$\left|\hat{p}_{y}(x)-p_{y}(x)\right|\leq{}$ $\underbrace{\left|\hat{c}_{y}(\hat{u})-c_{y}(u)\right|}_{\text{Copula error}}\prod_{i}\hat{f}_{i|y}(x_{i})$
$+c_{y}(u)\underbrace{\left|\prod_{i}\hat{f}_{i|y}(x_{i})-\prod_{i}f_{i|y}(x_{i})\right|}_{\text{Marginal product error}}$
Using the Portmanteau lemma [29], uniform convergence of $\hat{p}_{y}\to p_{y}$ implies:
$\arg\max_{y}\hat{p}_{y}(x)\to_{p}\arg\max_{y}p_{y}(x)$
completing the Bayes consistency argument. ∎
A.2 Proof of Theorem 3 (Convergence Rate)
The excess risk decomposes as:
Approximation Error
For copula densities in $W^{r}_{2}([0,1]^{d})$, [25] gives ReLU networks with $L=O(\log n)$ layers and $W=O(n^{d/(2r+d)})$ neurons per layer achieving:
$\inf_{\theta_{y}}\|NN_{y}(\cdot;\theta_{y})-c_{y}\|_{L^{2}}\leq Cn^{-\frac{r}{2r+d}}$
Estimation Error
Using Gaussian complexity bounds [farrell2021deep], the MLE satisfies with probability $1-\delta$:
$\mathbb{E}_{u}\left[\left|NN_{y}(u;\hat{\theta}_{y})-c_{y}(u)\right|^{2}\right]\leq C\left(n^{-\frac{2r}{2r+d}}+\frac{\log(1/\delta)}{n}\right)$
Marginal Contribution
Under Assumption 1(5), the product density term contributes:
$\left|\prod_{i=1}^{d}\hat{f}_{i|y}(x_{i})-\prod_{i=1}^{d}f_{i|y}(x_{i})\right|=O_{p}(n^{-\alpha})$
Composite Rate
Combining via Hölder’s inequality:
$E(n)\leq C\left(n^{-\frac{r}{2r+d}}+n^{-\alpha}+n^{-1/2}\right)$
The dominant term is $n^{-\frac{r}{2r+d}}$ since $\alpha>\frac{r}{2r+d}$ by assumption. ∎
Remark: The logarithmic factors in Theorem 3 arise from covering number bounds in [25] and union bounds over $K$ classes.