### v2

# Theoretical Foundations of the Deep Copula Classifier: A Generative Approach to Modeling Dependent Features

Abstract
Traditional classifiers often assume feature independence or rely on overly simplistic relationships, leading to poor performance in settings where real-world dependencies matter. In this paper, we introduce the Deep Copula Classifier (DCC), a generative model that separates the learning of each feature’s marginal distribution from the modeling of their joint dependence structure via neural network-parameterized copulas. For each class, lightweight neural networks are used to flexibly and adaptively capture feature interactions, making DCC particularly effective when classification is driven by complex dependencies. We establish that DCC converges to the Bayes-optimal classifier under standard conditions and provide explicit convergence rates of $O(n^{-r/(2r+d)})$ for $r$-smooth copula densities. Beyond theoretical guarantees, we outline several practical extensions, including high-dimensional scalability through vine and factor copula architectures, semi-supervised learning via entropy regularization, and online adaptation using streaming gradient methods. By unifying statistical rigor with the representational power of neural networks, DCC offers a mathematically grounded and interpretable framework for dependency-aware classification.
Keywords:Classification, Copula Theory, Deep Generative Models, Feature Dependencies, Bayes Consistency, Neural Copulas
MSC 2020 Classification: 62H30, 68T07, 62C12, 62G05
1 Introduction
Classification, that is the task of assigning labels based on features, is a cornerstone of machine learning. Classical methods like logistic regression ([10]), support vector machines ([6]), and naive Bayes ([21]) succeed in many scenarios but rely on restrictive assumptions: naive Bayes presumes feature independence, linear discriminant analysis requires Gaussian features with shared covariance, and even neural networks encode feature interactions opaquely through weight matrices rather than explicit probabilistic modeling.
Real-world features, however, interact through complex nonlinear dependencies. Financial indicators exhibit extreme co-movements invisible to correlation ([8]), while in medical diagnostics, biomarkers like blood pressure and cholesterol levels may exhibit nonlinear dependencies critical for accurate disease prediction. Ignoring these interdependencies degrades classifier performance. Copulas elegantly solve this by decoupling marginal distributions from dependence structures ([30]). Traditional parametric copulas (Gaussian, Clayton, Gumbel) and newer Archimedean variants ([1]) capture basic patterns but falter in high dimensions or with intricate dependencies. Neural copulas ([33, 35, 13]) overcome these limitations through adaptive deep learning.
We bridge this gap with the Deep Copula Classifier (DCC), which for each class models individual features via their marginal distributions and learns feature dependencies through neural network-parameterized copulasThis generative approach explicitly quantifies how features co-vary, enabling classifiers that harness dependency patterns. Our theoretical contributions are as follows:
  1. 1.
    Novel Framework: Integration of deep copulas into a class-conditional generative classifier
  2. 2.
    Consistency Proof: DCC converges to Bayes-optimality under standard conditions
  3. 3.
    Rate Quantification: $O(n^{-r/(2r+d)})$ excess risk for $r$-smooth copulas
  4. 4.
    Interpretability: Direct inspection of learned copula dependencies
  5. 5.
    Extensions Roadmap: High-dimensional, semi-supervised, and streaming adaptations
DCC combines copula theory’s rigor with neural networks’ flexibility, offering a principled path for dependency-aware classification. This foundation supports both theoretical analysis and practical implementation.
2 Related Work
2.1 Copula Theory and Foundations
The copula framework, formalized by Sklar’s theorem ([30]), enables the separation of marginal distributions from joint dependence structures. Modern treatments, such as ([24]) establish copulas as foundational tools in multivariate analysis. Significant progress has been made in several directions. Nonparametric estimation techniques, including kernel-based copula density estimators ([25]) and empirical bayes method ([20]), have enabled flexible inference without strong distributional assumptions. In high dimensions, vine copulas ([23]) offer a scalable strategy by factorizing complex dependencies into cascades of bivariate copulas. Additionally, the integration of copulas into generative modeling frameworks, such as copula-based variational inference ([31]) and optimal transport-based constructions ([5]), has expanded their utility in probabilistic learning.
2.2 Generative Classification Paradigms
Traditional generative classifiers, such as naive Bayes ([21]), prioritize tractability over accurate modeling of feature dependence, often assuming conditional independence. In contrast, modern deep generative approaches address this limitation through implicit and explicit mechanisms. Models like variational autoencoders (VAEs) ([16]) and generative adversarial networks (GANs) ([12]) capture feature interactions within nonlinear latent spaces, providing powerful implicit dependence structures. Others, such as neural graphical model hybrids ([28]), incorporate explicit conditional independence assumptions to combine interpretability with deep learning. The Deep Copula Classifier (DCC) diverges from both approaches by directly parameterizing feature dependencies through copulas while preserving full generative semantics, a principled middle ground between rigid parametric assumptions and black-box neural inference.
2.3 Neural Copula Architectures
Recent advances in neural dependence modeling provide key building blocks for flexible and expressive copula-based models. Universal parameterization has been achieved using deep Archimedean copulas ([18]) and transformer-based copula processes ([35]), allowing for rich representations of complex multivariate dependencies. For adaptive estimation, nonparametric vine copulas augmented with neural pair-copula blocks have shown promise in capturing localized interactions in high dimensions ([23]). Additionally, hybrid learning paradigms such as neural copula processes ([13]) effectively blend generative and discriminative objectives, offering a unified framework for robust prediction and inference. Unlike prior neural copula models (e.g., [18]), DCC integrates class-conditional copulas with explicit convergence guarantees, enabling interpretable dependency modeling tailored to classification.
Our work extends these foundations by developing:
  1. 1.
    Class-specific copula networks with theoretical consistency guarantees
  2. 2.
    Explicit finite-sample convergence rates under smoothness conditions
  3. 3.
    Architectural constraints ensuring valid density outputs
    This bridges the gap between empirical neural copula methods and rigorous statistical learning theory.
3 Theoretical Framework
Definition 1 (Copula).
A $d$-dimensional copula is a function $C:[0,1]^{d}\to[0,1]$ satisfying:
  1. 1.
    Groundedness: If any $u_{i}=0$, then
[…]
  2. 2.
    Uniform margins: For each $i=1,\dots,d$,
[…]
  3. 3.
    $d$-increasing: For every pair of points $\mathbf{a}=(a_{1},\dots,a_{d})$ and $\mathbf{b}=(b_{1},\dots,b_{d})$ with $0\leq a_{j}\leq b_{j}\leq 1$, the “volume” of the hyper-rectangle $[\mathbf{a},\mathbf{b}]$ must be nonnegative. Equivalently,
[…]
    where, for each $j=1,\dots,d$,
    $w_{j}(\epsilon_{j})\;=\;\begin{cases}a_{j},&\epsilon_{j}=0,\\ b_{j},&\epsilon_{j}=1.\end{cases}$
    Here $\lvert\boldsymbol{\epsilon}\rvert=\sum_{j=1}^{d}\epsilon_{j}$ is the number of ones in $\boldsymbol{\epsilon}$. This property ensures the copula assigns nonnegative probability to hyper-rectangles, preserving multivariate monotonicity.
Note: In the case $d=2$, this “2‐increasing” condition reduces to:
[…]
ensuring that $C$ assigns nonnegative “mass” to each axis‐aligned rectangle in $[0,1]^{2}$.
3.1 Deep Copula Classifier Definition
Our classifier models each class’s joint distribution through neural copula parameterization.
Definition 2 (Deep Copula Classifier).
Let $X=(X_{1},\dots,X_{d})$ be a feature vector and $Y\in\{1,\dots,K\}$ a discrete label. For each class $y$, we decompose the conditional density as:
$P(X\mid Y=y)=c_{y}(F_{1|y}(X_{1}),\dots,F_{d|y}(X_{d}))\cdot\prod_{i=1}^{d}f_{i|y}(X_{i}),$
where:
  • $F_{i|y}$ and $f_{i|y}$ are the class-conditional marginal CDF and PDF;
  • $c_{y}$ is a copula density implemented through a neural network $NN_{y}(\cdot;\theta_{y})$ with softplus output activation and implicit normalization.
The transformation $u_{i}=F_{i|y}(X_{i})$ maps features to the unit hypercube, where $NN_{y}$ learns dependence patterns. While practical implementations require explicit density normalization, our theoretical analysis assumes the network can approximate any smooth copula density through Hornik’s universal approximation theorem [14].
3.2 Consistency and Convergence
We establish fundamental guarantees under the following conditions:
Regularity Conditions: Fix a class label $y\in\{1,\dots,K\}$. We assume the following hold for every $y$:
  1. (1)
    Existence of a copula‐based factorization. The conditional distribution of $X=(X_{1},\dots,X_{d})$ given $\{Y=y\}$ admits a continuous joint density
    $f_{X\mid Y=y}(x_{1},\dots,x_{d})$
    with strictly increasing marginal CDFs
    $F_{i\mid y}(x_{i})\;=\;\mathbb{P}(X_{i}\leq x_{i}\mid Y=y),\quad i=1,\dots,d.$
    By Sklar’s theorem (Definition 1 and Sklar [30]), there is a unique copula function $C_{y}:[0,1]^{d}\to[0,1]$ whose density
    $c_{y}(u_{1},\dots,u_{d})\;=\;\frac{\partial^{d}}{\partial u_{1}\,\cdots\,\partial u_{d}}\,C_{y}(u_{1},\dots,u_{d})$
    satisfies, for all $(x_{1},\dots,x_{d})\in\mathbb{R}^{d}$,
    $f_{X\mid Y=y}(x_{1},\dots,x_{d})\;=\;c_{y}\bigl{(}F_{1\mid y}(x_{1}),\,\dots,\,F_{d\mid y}(x_{d})\bigr{)}\;\times\;\prod_{i=1}^{d}f_{i\mid y}(x_{i}),$
    where $f_{i\mid y}(x_{i})=\frac{d}{dx_{i}}\,F_{i\mid y}(x_{i})$ is the $i$th marginal density. In other words, the class‐conditional law factorizes exactly as “copula density $\times$ marginals.”
  2. (2)
    Marginal regularity. Each marginal CDF $F_{i\mid y}(x_{i})$ is continuous and strictly increasing on $\mathcal{X}_{i|y}$, and its density $f_{i\mid y}(x_{i})$ exists and satisfies
    $0<m\;\leq\;f_{i\mid y}(x_{i})\;\leq\;M<\infty\quad\text{for all }x_{i}\in\mathcal{X}_{i|y}\subseteq\mathbb{R},$
    for some constants $m,M>0$. This ensures the transformation $u_{i}=F_{i\mid y}(x_{i})$ is well‐defined and invertible.
  3. (3)
    Neural‐copula approximation capacity. For each class $y$, the family of copula networks $\{NN_{y}(\,\cdot\,;\theta)\}$ has width and depth growing at most polynomially in $n$. By Hornik’s universal approximation theorem ([14]), this guarantees that for any Sobolev‐smooth copula density $c_{y}\in W_{2}^{r}([0,1]^{d})$, there exist parameters $\theta_{y}^{*}$ such that
    $\sup_{u\in[0,1]^{d}}\bigl{\lvert}NN_{y}(u;\theta_{y}^{*})\;-\;c_{y}(u)\bigr{\rvert}\;=\;O\!\bigl{(}n^{-r/(2r+d)}\bigr{)},$
    up to logarithmic factors, provided the network’s width and depth are chosen as in Theorem 2.
  4. (4)
    i.i.d. sampling. The training pairs $\{(X^{(i)},Y^{(i)})\}_{i=1}^{n}$ are independent and identically distributed samples from the true joint law $P_{X,Y}$. Thus we can invoke standard concentration and empirical‐process arguments (e.g. [32]).
  5. (5)
    Marginal‐estimator convergence. For each class $y$ and each feature index $i$, let $\widehat{F}_{i\mid y}$ be a nonparametric estimator (e.g. kernel‐smoothed empirical CDF) of $F_{i\mid y}$. We assume
    $\sup_{x\in\mathbb{R}}\Bigl{|}\widehat{F}_{i\mid y}(x)\;-\;F_{i\mid y}(x)\Bigr{|}\;=\;O_{p}\bigl{(}n^{-\alpha}\bigr{)}\quad\text{for some }\alpha>\frac{r}{\,2r+d\,},$
    In particular, $\alpha>r/(2r+d)$ ensures that marginal estimation errors do not dominate the copula approximation error, preserving overall convergence.
In the above and throughout, $n$ denotes the total number of i.i.d. training samples and $r$ is the Sobolev order of the true copula density $c_{y}\in W_{2}^{r}([0,1]^{d})$.
Theorem 1 (Consistency).
Let
$\widehat{Y}_{n}\;=\;\widehat{Y}_{n}(X)$
be the classifier produced by the Deep Copula Classifier (DCC) when all “hat” quantities (such as $\widehat{F}_{i\mid y}$, $\widehat{f}_{i\mid y}$, $\widehat{c}_{y}$, etc.) are estimated from $n$ i.i.d. training samples. In particular, $\widehat{Y}_{n}$ depends on $n$ because it uses the estimated class–conditional densities
$\widehat{p}_{y}^{(n)}(x)\;=\;\widehat{c}_{y}^{(n)}\bigl{(}\widehat{F}_{1\mid y}^{(n)}(x_{1}),\,\dots,\,\widehat{F}_{d\mid y}^{(n)}(x_{d})\bigr{)}\;\times\;\prod_{i=1}^{d}\widehat{f}_{i\mid y}^{(n)}(x_{i}),$
as outputs of Algorithm 2 trained on those $n$ samples. Define the Bayes‐optimal decision rule
$Y^{*}(x)\;=\;\arg\max_{y\in\{1,\dots,K\}}p_{y}(x),$
where $p_{y}(x)$ is the true class‐conditional density. Under Assumption 3.2 (with network capacity growing appropriately as $n\to\infty$), we have
$\lim_{n\to\infty}\;\mathbb{P}\bigl{(}\widehat{Y}_{n}(X)\neq Y\mid X\bigr{)}\;=\;\mathbb{P}\bigl{(}Y^{*}(X)\neq Y\mid X\bigr{)}.$
In other words, the sequence of data‐driven classifiers $\{\widehat{Y}_{n}\}_{n\geq 1}$ is Bayes‐consistent.
Theorem 2 (Convergence Rate).
Let
$\widehat{Y}_{n}\;=\;\widehat{Y}_{n}(X)$
be the classifier produced by DCC when all “hat” quantities are estimated from $n$ i.i.d. training samples (so in particular $\widehat{p}_{y}^{(n)}(x)$ depends on $n$). Define the excess risk
$E(n)\;=\;\mathbb{P}\bigl{(}\widehat{Y}_{n}(X)\neq Y\bigr{)}\;-\;\mathbb{P}\bigl{(}Y^{*}(X)\neq Y\bigr{)},\qquad Y^{*}(x)\;=\;\arg\max_{y}p_{y}(x).$
Suppose that:
  1. 1.
    For each class $y$, the true copula density $c_{y}$ lies in the Sobolev space $W_{2}^{r}([0,1]^{d})$ of order $r>d/2$.
  2. 2.
    Assumption 3.2 holds, in particular so that the marginal‐CDF estimators satisfy $\sup_{x}|\widehat{F}_{i\mid y}(x)-F_{i\mid y}(x)|=O_{p}(n^{-\alpha})$ with
    $\alpha\;>\;\frac{r}{2r+d}.$
Then, as $n\to\infty$, the excess risk satisfies
$E(n)\;=\;\mathbb{P}\bigl{(}\widehat{Y}_{n}(X)\neq Y\bigr{)}\;-\;\mathbb{P}\bigl{(}Y^{*}(X)\neq Y\bigr{)}\;=\;O_{p}\!\Bigl{(}n^{-\,\tfrac{r}{2r+d}}+n^{-\alpha}\Bigr{)}\;=\;O_{p}\!\Bigl{(}n^{-\,\tfrac{r}{2r+d}}\Bigr{)},$
neglecting any logarithmic factors. In other words, the excess risk of the DCC classifier decays at the rate $n^{-\,r/(2r+d)}$.
For detailed proofs, See Appendix Appendix A: Detailed Proofs.
4 Algorithm and Implementation
4.1 Training Procedure
The DCC training process decomposes into marginal estimation and dependence learning, as formalized in Algorithm 1.
Algorithm 1 Theoretical Training Protocol for DCC
1:Labeled sample $\{(x^{(i)},\,y^{(i)})\}_{i=1}^{n}\stackrel{{\scriptstyle iid}}{{\sim}}P_{X,Y}$
2:Estimated marginals $\{\hat{F}_{j|y},\,\hat{f}_{j|y}\}$ and network parameters $\{\hat{\theta}_{y}\}_{y=1}^{K}$
3:
4:Phase 1: Consistent Marginal Estimation
5:for each class $y\in\{1,\dots,K\}$ do
6:     for each feature $j\in\{1,\dots,d\}$ do
7:         Estimate $\hat{F}_{j|y}$ via a smoothed empirical CDF with bandwidth $h=o(n^{-1/2})$.
8:         Derive $\hat{f}_{j|y}\;=\;\tfrac{d}{dx}\,\hat{F}_{j|y}$ using kernel density estimation.
9:     end for
10:end for
11:
12:Phase 2: Neural Copula Learning
13:for each class $y\in\{1,\dots,K\}$ do
14:     Transform data: $u_{j}^{(i)}\leftarrow\hat{F}_{j|y}\bigl{(}x_{j}^{(i)}\bigr{)}$ for all $i$ such that $y^{(i)}=y$.
15:     Define neural copula network $NN_{y}$ (architecture in Section 4.3).
16:     Solve
$\hat{\theta}_{y}\;=\;\arg\max_{\theta}\;\frac{1}{\,n_{y}\,}\sum_{i:\,y^{(i)}=y}\log\Bigl{(}NN_{y}\bigl{(}u^{(i)};\theta\bigr{)}\Bigr{)},$
where $n_{y}=\bigl{\lvert}\{\,i:y^{(i)}=y\}\bigr{\rvert}$.
17:end for
Note: In practice, $h$ can be selected via cross-validation or Silverman’s rule ([29]) to balance bias-variance.
4.2 Prediction Procedure
Classification follows the Bayes decision rule under the estimated densities (Algorithm 2).
Algorithm 2 Theoretical Prediction Mechanism
1:Test point $x^{*}$, trained DCC model
2:Predicted class $\hat{y}$
3:
4:for each class $y\in\{1,\dots,K\}$ do
5:     Transform: $u_{j}^{*}\leftarrow\hat{F}_{j|y}\bigl{(}x_{j}^{*}\bigr{)}$ for $j=1,\dots,d$.
6:     Evaluate copula density: $\hat{c}_{y}\bigl{(}u^{*}\bigr{)}\leftarrow NN_{y}\bigl{(}u^{*};\hat{\theta}_{y}\bigr{)}$.
7:     Compute class‐conditional likelihood: $L_{y}\leftarrow\hat{c}_{y}\bigl{(}u^{*}\bigr{)}\;\times\;\prod_{j=1}^{d}\hat{f}_{j|y}\bigl{(}x_{j}^{*}\bigr{)}$.
8:     Incorporate prior: $\tilde{L}_{y}\leftarrow L_{y}\;\times\;\hat{\pi}_{y}$, where $\hat{\pi}_{y}=\tfrac{n_{y}}{n}$.
9:end for
10:return $\hat{y}=\arg\max_{y\in\{1,\dots,K\}}\tilde{L}_{y}$
4.3 Neural Network Architecture
The copula networks $NN_{y}$ require architectural constraints to ensure theoretical validity:
[…]
  • Hidden Layers: Depth $L=\lceil\log_{2}n\rceil$ with width $W=O(n^{d/(2r+d)})$ for $r$-smooth copulas
  • Activation: ReLU for hidden layers, Softplus output activation for positivity
  • Regularization: Implicit via spectral normalization ([22]) to control Rademacher complexity
Note: While theoretically achievable through Barron space approximation ([2]), practical implementations enforce the integral constraint via normalization techniques like softmax over a grid or Monte Carlo integration.
5 Conclusion and Future Work
Our theoretical analysis establishes the Deep Copula Classifier (DCC) as a Bayes-consistent method that explicitly models feature dependencies through neural copula parameterization. By separating marginal learning from dependence estimation, DCC provides an interpretable framework where practitioners can inspect both individual feature behaviors (through $\hat{F}_{j|y}$) and interaction patterns (via $NN_{y}$). The $O(n^{-r/(2r+d)})$ convergence rate for Sobolev-class copulas demonstrates adaptivity to intrinsic data complexity.
The following avenues for theoretical extension emerge:
High-Dimensional Scaling: The curse of dimensionality in $[0,1]^{d}$ copula estimation poses significant challenges. To address this, one promising direction involves vine copula factorizations ([23]), which break down high-dimensional dependencies into tree-structured graphs that are easier to estimate. Additionally, low-rank tensor decompositions of the weight matrices within $NN_{y}$ may reduce parameter complexity while preserving expressive power. Another line of work would analyze the tradeoff between sparsity and sample size in regimes where $d\gg n$, providing theoretical insights into when DCC remains statistically reliable.
Semi-Supervised Learning: Given its generative nature, DCC naturally lends itself to semi-supervised extensions under the cluster assumption ([4]). This opens the door to convergence analyses for entropy-regularized objectives that leverage both labeled and unlabeled data. In particular, DCC may benefit when marginal distributions $P(X)$ carry information about class boundaries, leading to provable gains in label efficiency. Furthermore, a rigorous comparison between DCC and discriminative semi-supervised methods could clarify the advantages of modeling dependencies explicitly.
Online Adaptation: In streaming data settings, DCC can be extended to incorporate online learning mechanisms. One direction is to establish regret bounds for sequential updates to the copula density via mirror descent ([15]). Beyond regret analysis, theoretical guarantees will be needed to understand how DCC behaves under class-incremental scenarios, especially regarding catastrophic forgetting. Additionally, DCC provides a unique opportunity to model and detect concept drift through shifts in learned dependence structures over time.
Empirical-Theoretical Synergy: Although our focus has been theoretical, empirical observations raise questions that demand mathematical exploration. For example, DCC may exhibit phase transitions in classification performance as the strength of feature dependencies varies, phenomena that could be characterized through theoretical thresholds. Another promising direction involves deriving information-theoretic lower bounds for copula-based classification, which would help identify when dependency modeling provides fundamental advantages over independence-based models.
The DCC framework opens new questions about balancing parametric and nonparametric components in hybrid models. By maintaining explicit probability models rather than black-box representations, it enables precise characterization of when and how feature dependencies impact classification, a crucial step toward explainable AI for high-stakes decisions.
Statements and Declarations
Funding
The authors did not receive any funding for this research.
Competing Interests
The authors have no relevant financial or non-financial interests to disclose.
Appendix A: Detailed Proofs
A.1 Proof of Theorem 1 (Consistency)
We prove that, under Assumption 3.2, the Deep Copula Classifier (DCC) converges to the Bayes‐optimal classifier as $n\to\infty$. Recall that every “hat” quantity ($\widehat{F}_{i\mid y},\,\widehat{f}_{i\mid y},\,\widehat{c}_{y},\,\widehat{p}_{y},\,\widehat{Y}$, etc.) depends on the total sample size $n$ because it is estimated from $n$ i.i.d. training pairs.
Step 1: Marginal Convergence
By Assumption 3.2(5), each marginal CDF estimator satisfies
$\sup_{x\in\mathbb{R}}\bigl{|}\widehat{F}_{i\mid y}(x)-F_{i\mid y}(x)\bigr{|}\;=\;O_{p}\bigl{(}n^{-\alpha}\bigr{)},\quad\alpha>\frac{r}{2r+d}.$
Because Assumption 3.2(2) ensures
$f_{i\mid y}(x)\;\geq\;m>0\quad\text{for all }x,$
each true marginal CDF $F_{i\mid y}$ is strictly increasing with derivative at least $m$. Hence its inverse $F_{i\mid y}^{-1}:[0,1]\to\mathbb{R}$ is Lipschitz with constant $1/m$. Concretely, for any $u,v\in[0,1]$:
$\bigl{|}\,F_{i\mid y}^{-1}(u)-F_{i\mid y}^{-1}(v)\bigr{|}\;\leq\;\tfrac{1}{m}\,\lvert u-v\rvert.$
Define, for each training point $X_{i}$,
$u_{i}=F_{i\mid y}(X_{i}),\qquad\widehat{u}_{i}=\widehat{F}_{i\mid y}(X_{i}).$
Then
$\lvert\widehat{u}_{i}-u_{i}\rvert\;=\;\bigl{|}\widehat{F}_{i\mid y}(X_{i})-F_{i\mid y}(X_{i})\bigr{|}\;=\;O_{p}\bigl{(}n^{-\alpha}\bigr{)}.$
Therefore the vector of pseudo‐observations $\widehat{\mathbf{u}}=(\widehat{u}_{1},\dots,\widehat{u}_{d})$ satisfies
$\bigl{\|}\widehat{\mathbf{u}}-\mathbf{u}\bigr{\|}_{\infty}\;=\;\max_{1\leq i\leq d}\lvert\widehat{u}_{i}-u_{i}\rvert\;=\;O_{p}\bigl{(}n^{-\alpha}\bigr{)}.$
Step 2: Neural Copula Approximation
Assumption 3.2(3) guarantees that our chosen class of neural networks can approximate any Sobolev‐smooth copula density. In particular, since the true class‐conditional copula $c_{y}\in W_{2}^{r}([0,1]^{d})$ with $r>d/2$, [27] shows there exists a ReLU network of depth $O(\log n)$ and width $O\bigl{(}n^{d/(2r+d)}\bigr{)}$ whose parameters $\theta_{y}^{*}$ satisfy
$\sup_{u\in[0,1]^{d}}\bigl{|}\,NN_{y}(u;\theta_{y}^{*})-c_{y}(u)\bigr{|}\;=\;O\!\bigl{(}n^{-\,r/(2r+d)}\bigr{)}.$
Let $\widehat{c}_{y}(u)=NN_{y}(u;\widehat{\theta}_{y})$ denote the network after training on $n_{y}$ samples. By standard Rademacher‐complexity‐based MLE arguments ([17]), one shows
$\mathbb{E}_{u\sim\mathrm{Unif}[0,1]^{d}}\bigl{[}\bigl{|}\,NN_{y}(u;\widehat{\theta}_{y})-c_{y}(u)\bigr{|}^{2}\bigr{]}\;=\;O_{p}\!\Bigl{(}n^{-\,\tfrac{2r}{2r+d}}\;+\;n^{-1}\Bigr{)}.$
Dropping the $O(n^{-1})$ term and applying Sobolev embedding (which upgrades $L^{2}$-approximation to $L^{\infty}$ up to logs) yields
$\sup_{u\in[0,1]^{d}}\bigl{|}\;\widehat{c}_{y}(u)-c_{y}(u)\bigr{|}\;=\;O_{p}\!\bigl{(}n^{-\,r/(2r+d)}\bigr{)}.$
Step 3: Decision Rule Stability
For any test input $x=(x_{1},\dots,x_{d})$, define
$p_{y}(x)\;=\;c_{y}\bigl{(}F_{1\mid y}(x_{1}),\,\dots,\,F_{d\mid y}(x_{d})\bigr{)}\;\times\;\prod_{i=1}^{d}f_{i\mid y}(x_{i}),$
and its estimate
$\widehat{p}_{y}(x)\;=\;\widehat{c}_{y}\bigl{(}\widehat{\mathbf{u}}\bigr{)}\;\times\;\prod_{i=1}^{d}\widehat{f}_{i\mid y}(x_{i}),\quad\text{where }\widehat{\mathbf{u}}=\bigl{(}\widehat{u}_{1},\dots,\widehat{u}_{d}\bigr{)}.$
We bound
$\bigl{|}\widehat{p}_{y}(x)-p_{y}(x)\bigr{|}$ $=\Bigl{|}\,\widehat{c}_{y}(\widehat{\mathbf{u}})\,\prod_{i=1}^{d}\widehat{f}_{i\mid y}(x_{i})\;-\;c_{y}(\mathbf{u})\,\prod_{i=1}^{d}f_{i\mid y}(x_{i})\Bigr{|}$
$\leq\underbrace{\bigl{|}\widehat{c}_{y}(\widehat{\mathbf{u}})-c_{y}(\mathbf{u})\bigr{|}\;\prod_{i=1}^{d}\widehat{f}_{i\mid y}(x_{i})}_{\text{(A) Copula‐error}}\;+\;\underbrace{\bigl{|}c_{y}(\mathbf{u})\bigr{|}\;\bigl{|}\prod_{i=1}^{d}\widehat{f}_{i\mid y}(x_{i})-\prod_{i=1}^{d}f_{i\mid y}(x_{i})\bigr{|}}_{\text{(B) Marginal‐error}}.$
(A) Copula‐error. Observe
$\bigl{|}\widehat{c}_{y}(\widehat{\mathbf{u}})-c_{y}(\mathbf{u})\bigr{|}\;\leq\;\bigl{|}\widehat{c}_{y}(\widehat{\mathbf{u}})-c_{y}(\widehat{\mathbf{u}})\bigr{|}\;+\;\bigl{|}c_{y}(\widehat{\mathbf{u}})-c_{y}(\mathbf{u})\bigr{|}.$
1. From Step 2, $\sup_{u}\lvert\widehat{c}_{y}(u)-c_{y}(u)\rvert=O_{p}\!\bigl{(}n^{-\,r/(2r+d)}\bigr{)}.$ 2. Since $c_{y}\in W_{2}^{r}$ with $r>d/2$, Sobolev embedding implies $c_{y}$ is Lipschitz on $[0,1]^{d}$. Let $L_{c}$ denote its Lipschitz constant. Then
$\bigl{|}\,c_{y}(\widehat{\mathbf{u}})-c_{y}(\mathbf{u})\bigr{|}\;\leq\;L_{c}\,\bigl{\|}\widehat{\mathbf{u}}-\mathbf{u}\bigr{\|}_{\infty}\;=\;O_{p}\bigl{(}n^{-\alpha}\bigr{)}\quad(\text{from Step 1}).$
Hence
$\sup_{x}\bigl{|}\widehat{c}_{y}(\widehat{\mathbf{u}})-c_{y}(\mathbf{u})\bigr{|}\;=\;O_{p}\!\bigl{(}n^{-\,\tfrac{r}{2r+d}}+n^{-\alpha}\bigr{)}.$
Since each $\widehat{f}_{i\mid y}(x_{i})$ is $O_{p}(1)$ uniformly (Assumption 3.2(2)), it follows that
$\sup_{x}\Bigl{|}\widehat{c}_{y}(\widehat{\mathbf{u}})-c_{y}(\mathbf{u})\Bigr{|}\,\prod_{i}\widehat{f}_{i\mid y}(x_{i})\;=\;O_{p}\!\bigl{(}n^{-\,\tfrac{r}{2r+d}}+n^{-\alpha}\bigr{)}.$
(B) Marginal‐error. Write $g_{i}(x_{i})=f_{i\mid y}(x_{i})$ and $\widehat{g}_{i}(x_{i})=\widehat{f}_{i\mid y}(x_{i})$. Then
$\prod_{i=1}^{d}\widehat{g}_{i}(x_{i})-\prod_{i=1}^{d}g_{i}(x_{i})\;=\;\sum_{j=1}^{d}\bigl{(}\widehat{g}_{j}(x_{j})-g_{j}(x_{j})\bigr{)}\;\prod_{i<j}\widehat{g}_{i}(x_{i})\;\prod_{i>j}g_{i}(x_{i}).$
By Step 1 and standard kernel‐density theory ([11]), $\sup_{x_{i}}\lvert\widehat{g}_{i}(x_{i})-g_{i}(x_{i})\rvert=O_{p}(n^{-\alpha})$. Since each $\widehat{g}_{i}$ and $g_{i}$ is bounded in probability by some constant $M$, each summand is
$O_{p}(n^{-\alpha})\times M^{\,d-1}=O_{p}(n^{-\alpha}).$
Because there are $d$ terms,
$\sup_{x}\bigl{|}\prod_{i=1}^{d}\widehat{g}_{i}(x_{i})-\prod_{i=1}^{d}g_{i}(x_{i})\bigr{|}\;=\;O_{p}\bigl{(}n^{-\alpha}\bigr{)}.$
Since $c_{y}(\mathbf{u})$ is bounded by Sobolev embedding, it follows that
$\sup_{x}\Bigl{|}\,c_{y}(\mathbf{u})\,\bigl{(}\prod_{i}\widehat{f}_{i\mid y}(x_{i})-\prod_{i}f_{i\mid y}(x_{i})\bigr{)}\Bigr{|}\;=\;O_{p}\bigl{(}n^{-\alpha}\bigr{)}.$
Putting (A) and (B) Together
Combining the bounds for (A) and (B) gives
$\sup_{x}\bigl{|}\,\widehat{p}_{y}(x)-p_{y}(x)\bigr{|}\;=\;O_{p}\!\bigl{(}n^{-\,\tfrac{r}{2r+d}}+n^{-\alpha}\bigr{)}.$
Since $\alpha>\tfrac{r}{2r+d}$ by Assumption 3.2(5), the term $n^{-\alpha}$ is asymptotically smaller, so
$\sup_{x}\bigl{|}\,\widehat{p}_{y}(x)-p_{y}(x)\bigr{|}\;=\;O_{p}\!\bigl{(}n^{-\,\tfrac{r}{2r+d}}\bigr{)}.$
Uniform convergence $\widehat{p}_{y}\to p_{y}$ for all $y=1,\dots,K$ implies
[…]
by a standard argmax‐continuity argument (e.g. Portmanteau lemma ([32]) plus union bound over the finitely many classes). Therefore $\widehat{Y}$ converges in probability to the Bayes‐optimal decision $Y^{*}$. This completes the proof of Theorem 1.
A.2 Proof of Theorem 2 (Convergence Rate)
We now show that, under the same regularity assumptions plus $c_{y}\in W_{2}^{r}([0,1]^{d})$, the excess classification risk decays at rate $n^{-\,r/(2r+d)}$.
Define
$E(n)\;=\;\mathbb{P}\bigl{(}\widehat{Y}_{n}(X)\neq Y\bigr{)}\;-\;\mathbb{P}\bigl{(}Y^{*}(X)\neq Y\bigr{)},\quad Y^{*}(x)\;=\;\arg\max_{y}p_{y}(x).$
By [7],
$E(n)\;\leq\;\sum_{y=1}^{K}\mathbb{E}_{X}\bigl{[}\bigl{|}\widehat{p}_{y}(X)-p_{y}(X)\bigr{|}\bigr{]}\;=\;\sum_{y=1}^{K}\bigl{\|}\widehat{p}_{y}-p_{y}\bigr{\|}_{L^{1}}.$
Hence it suffices to bound $\|\widehat{p}_{y}-p_{y}\|_{L^{1}}$ for each $y$. We decompose this into three parts, exactly as in Appendix A.1:
  1. 1.
    Copula‐Approximation Error: By [27], there exists a ReLU network of depth $O(\log n)$, width $O\bigl{(}n^{d/(2r+d)}\bigr{)}$, and parameters $\theta_{y}^{*}$ such that
    $\inf_{\theta}\bigl{\|}\,NN_{y}(\cdot;\theta)-c_{y}\bigr{\|}_{L^{2}([0,1]^{d})}\;\leq\;C\,n^{-\,r/(2r+d)}.$
    Sobolev embedding (since $r>d/2$) implies
    $\sup_{u\in[0,1]^{d}}\bigl{|}\,NN_{y}(u;\theta_{y}^{*})-c_{y}(u)\bigr{|}\;=\;O\!\bigl{(}n^{-\,r/(2r+d)}\bigr{)}.$
    Consequently,
    $\bigl{\|}\,NN_{y}(\cdot;\theta_{y}^{*})\,\prod_{i}f_{i\mid y}\;-\;c_{y}\,\prod_{i}f_{i\mid y}\bigr{\|}_{L^{1}}\;=\;O\!\bigl{(}n^{-\,r/(2r+d)}\bigr{)},\quad\Bigl{(}\int\prod_{i}f_{i\mid y}=1\Bigr{)}.$
  2. 2.
    Network‐Estimation Error. Let
    $\begin{split}\mathcal{F}_{y,n}&=\bigl{\{}\,u\mapsto NN_{y}(u;\theta)\;:\;\theta\text{ ranges over all ReLU‐networks of depth }L_{n}=O(\log n)\\ &\qquad\quad\text{and width }W_{n}=O\bigl{(}n^{d/(2r+d)}\bigr{)}\bigr{\}}.\end{split}$
    We train $\widehat{\theta}_{y}$ by maximizing the log‐likelihood of $\widehat{c}_{y}(u)=NN_{y}(u;\widehat{\theta}_{y})$ on the $n_{y}$ “pseudo‐observations” $\{\,u^{(i)}=\widehat{\mathbf{u}}^{(i)}\mid y^{(i)}=y\,\}$, each distributed roughly like $\mathrm{Unif}[0,1]^{d}$.
    In [9], it is shown that if a function class $\mathcal{F}$ has Gaussian‐complexity (or Rademacher‐complexity) bounded by $\mathfrak{R}(n)$, then the empirical‐risk minimizer $\widehat{f}$ over $\mathcal{F}$ satisfies with probability at least $1-\delta$:
    $\mathbb{E}_{u\sim\mathrm{Unif}[0,1]^{d}}\bigl{[}\bigl{|}\,\widehat{f}(u)-c_{y}(u)\bigr{|}^{2}\bigr{]}\;\leq\;\inf_{f\in\mathcal{F}}\mathbb{E}_{u}\bigl{[}\bigl{|}\,f(u)-c_{y}(u)\bigr{|}^{2}\bigr{]}\;+\;O\!\Bigl{(}\mathfrak{R}(n)\Bigr{)}\;+\;O\!\bigl{(}n^{-1}\bigr{)}.$
    In our setting:
[…]
    • The Gaussian complexity $\mathfrak{R}(n)$ of $\mathcal{F}_{y,n}$ can be bounded by
      $\mathfrak{R}(n)\;\approx\;O\!\Bigl{(}\sqrt{\tfrac{L_{n}\,W_{n}}{n}}\Bigr{)}\;=\;O\!\Bigl{(}\sqrt{\tfrac{(\log n)\,n^{\,d/(2r+d)}}{n}}\Bigr{)}\;=\;O\!\Bigl{(}n^{-\,\tfrac{r}{2r+d}}\Bigr{)},$
      up to logarithmic factors. Concretely, standard covering‐number or Rademacher‐complexity bounds for ReLU networks ( [3]; [17]) show that
      $\mathfrak{R}(n)\;=\;O\!\Bigl{(}\sqrt{\tfrac{LW}{n}}\Bigr{)}\quad\text{whenever}\quad L=O(\log n),\ W=O\bigl{(}n^{d/(2r+d)}\bigr{)}.$
      Hence $\mathfrak{R}(n)=O\bigl{(}n^{-\,r/(2r+d)}\bigr{)}$.
[…]
    $\mathbb{E}_{u}\bigl{[}\bigl{|}\,NN_{y}(u;\widehat{\theta}_{y})-c_{y}(u)\bigr{|}^{2}\bigr{]}\;=\;O\!\Bigl{(}n^{-\,\tfrac{2r}{2r+d}}\Bigr{)}\;+\;O\!\Bigl{(}n^{-\,\tfrac{r}{2r+d}}\Bigr{)}\;+\;O\!\bigl{(}n^{-1}\bigr{)}.$
    Since $r>d/2$, the largest of these exponents is $\tfrac{2r}{2r+d}>\tfrac{r}{2r+d}$. In big‐$O$ notation, one thus writes
    $\mathbb{E}_{u}\bigl{[}\bigl{|}\,NN_{y}(u;\widehat{\theta}_{y})-c_{y}(u)\bigr{|}^{2}\bigr{]}\;=\;O_{p}\!\Bigl{(}n^{-\,\tfrac{2r}{2r+d}}+n^{-1}\Bigr{)}.$
    Finally, because $c_{y}\in W_{2}^{r}([0,1]^{d})$ with $r>d/2$, Sobolev‐embedding guarantees an $L^{2}\to L^{\infty}$ upgrade (up to logs):
    $\sup_{u\in[0,1]^{d}}\bigl{|}\widehat{c}_{y}(u)-c_{y}(u)\bigr{|}\;=\;O_{p}\!\bigl{(}n^{-\,r/(2r+d)}\bigr{)}.$
    Hence overall
    $\bigl{\|}\,NN_{y}(\cdot;\widehat{\theta}_{y})\,\prod_{i}f_{i\mid y}\;-\;c_{y}\,\prod_{i}f_{i\mid y}\bigr{\|}_{L^{1}}\;=\;O_{p}\!\bigl{(}n^{-\,r/(2r+d)}\bigr{)}.$
  3. 3.
    Marginal‐Estimation Error: Under Assumption 3.2(5),
    $\sup_{x_{i}}\bigl{|}\widehat{F}_{i\mid y}(x_{i})-F_{i\mid y}(x_{i})\bigr{|}\;=\;O_{p}\bigl{(}n^{-\alpha}\bigr{)},\quad\alpha>\frac{r}{2r+d}.$
    By standard kernel‐density theory ([11]), $\sup_{x_{i}}\lvert\widehat{f}_{i\mid y}(x_{i})-f_{i\mid y}(x_{i})\rvert=O_{p}(n^{-\alpha})$. Since each $f_{i\mid y}$ is bounded away from zero and infinity, a telescoping product‐difference argument shows
    $\sup_{x}\bigl{|}\prod_{i=1}^{d}\widehat{f}_{i\mid y}(x_{i})-\prod_{i=1}^{d}f_{i\mid y}(x_{i})\bigr{|}\;=\;O_{p}\bigl{(}n^{-\alpha}\bigr{)}.$
    Moreover, $c_{y}$ is bounded by Sobolev embedding, so
    $\bigl{\|}\,c_{y}\,\bigl{(}\prod_{i}\widehat{f}_{i\mid y}-\prod_{i}f_{i\mid y}\bigr{)}\bigr{\|}_{L^{1}}\;=\;O_{p}\bigl{(}n^{-\alpha}\bigr{)}.$
Combining 1,2 and 3, we conclude for each class $y$:
$\bigl{\|}\widehat{p}_{y}-p_{y}\bigr{\|}_{L^{1}}\;=\;O_{p}\!\bigl{(}n^{-\,r/(2r+d)}+n^{-\alpha}\bigr{)}\;=\;O_{p}\!\bigl{(}n^{-\,r/(2r+d)}\bigr{)},$
since $\alpha>r/(2r+d)$. Summing over $y=1,\dots,K$ (a constant factor) yields
$E(n)\;=\;\sum_{y=1}^{K}\|\widehat{p}_{y}-p_{y}\|_{L^{1}}\;=\;O_{p}\!\bigl{(}n^{-\,r/(2r+d)}\bigr{)}.$
This completes the proof of Theorem 2.
Remark. Any extra $\log n$ or $\log(1/\delta)$ factors from covering‐number or union‐bound arguments are absorbed into the $O_{p}(\cdot)$ notation and do not affect the leading exponent $r/(2r+d)$.